# Majority-Logic Decoding of Binary Locally Recoverable Codes: A Probabilistic Analysis

### v1

@@ IV Simulations
To validate the theoretical bounds derived for MLD over the BSC, we performed simulations analyzing both BER and BLER. The simulations utilized a fixed locality parameter of $r=4$ while varying the availability $t(n)$ across different growth regimes relative to the block length $n$. The plotted numerical values are each averaged over $5\times 10^{6}$ runs.
@@ IV-A Bit Decoding Failure Probability
  • Linear and Polylogarithmic Regimes: In cases where $t$ grows linearly or as $(\log_{2}n)^{2}$, the empirical failure probability decays rapidly to zero as $n$ increases, confirming that MLD successfully corrects errors when the number of recovery sets is sufficiently large. Notably, the empirical performance is consistently superior to the theoretical upper bounds, though the gap narrows as $n$ grows. The tightness of the bound is directly related to the growth rate of $t(n)$; for the linear case ($t=n/4$), the theoretical bound closely approximates the empirical probability for $n\geq 2^{12}$, whereas it provides a looser estimate in the slower-growth regimes (e.g., $t=\sqrt{\log n}$). Note that as mentioned in Remark 1, although being exponentially tight, the theoretical upper bound can be further refined by incorporating a multiplicative pre-factor.
  • Sub-logarithmic Regime: Conversely, for the slowly growing availability $t=\sqrt{\log_{2}n}$, the decoding failure probability does not vanish. Instead, it saturates at a high error floor (approximately 0.2 to 0.4), indicating that when availability grows slower than logarithmic factors, the majority vote may fail to correct errors reliably.
@@ IV-B Block Decoding Failure Probability
Figure 2 focuses on the upper bound of the BLER, derived via the union bound on the Bit Error Rate (BER) (i.e., $P_{\mathrm{fail,block}}\leq nP_{\mathrm{fail,bit}}$) for a BSC with $p_{f}=0.13$. Since this BLER is derived from a union bound, the resulting values may exceed both the true BLER and even $1$, which is reflected in the vertical axis of the plot. We do not report empirical BLER values, as their computation becomes infeasible for large blocklengths: verifying correct block decoding requires checking the correctness of all symbols, each of which involves an increasing number of recovery sets as $n$ grows. This figure provides a fine-grained analysis of the polylogarithmic regime by varying the exponent of the logarithmic growth, specifically $t=(\log_{2}n)^{\alpha}$ for $\alpha\in\{1.8,1.9,2.05\}$. Figure 2 highlights the sensitivity of block-level decoding success to the exponent $\alpha$:
  • Failure Regime ($\alpha=1.8$): For $t=(\log_{2}n)^{1.8}$, the bound on BLER initially rises and remains high, failing to converge to 0 until $n\geq 2^{23}$. This suggests that this growth rate is insufficient to ensure vanishing BLER at practical block lengths.
  • Success Regimes ($\alpha=1.9,2.05$): As the exponent increases, the failure probability transitions into a clear decay phase, vanishing significantly faster than in the previous case. This confirms that for block-level reliability—which requires every single bit to be correct—the availability must scale with a sufficiently high power of $\log n$ to counteract the cumulative bit error probability resulting from the increasing block length.
@@ Acknowledgment
This work was supported in part by the NSF-BSF grant FET212026. The authors thank Alexander Barg for providing them with an original manuscript in Russian and its English translation.

### v2

@@ IV Simulations
To validate the theoretical bounds derived for MLD over the BSC, we performed simulations analyzing both BER and BLER. The simulations utilized a fixed locality parameter of $r=4$ while varying the availability $t(n)$ across different growth regimes relative to the block length $n$. We intentionally choose a relatively large crossover probability $p_{f}$ to demonstrate the robustness of MLD against high-weight error patterns when sufficient availability is provided. The plotted numerical values are each averaged over $5\times 10^{6}$ runs.
@@ IV-A Bit Decoding Failure Probability
  • Linear and Polylogarithmic Regimes: In cases where $t$ grows linearly or as $(\log_{2}n)^{2}$, the empirical failure probability decays rapidly to zero as $n$ increases, confirming that MLD successfully corrects errors when the number of recovery sets is sufficiently large. Notably, the empirical performance is consistently superior to the theoretical upper bounds, though the gap narrows as $n$ grows. The tightness of the bound is directly related to the growth rate of $t(n)$; for the linear case ($t=n/4$), the theoretical bound closely approximates the empirical probability for $n\geq 2^{12}$, whereas it provides a much looser estimate in the slower-growth regimes (e.g., $t=\sqrt{\log n}$). Note that as mentioned in Remark 1, although being exponentially tight, the theoretical upper bound can be further refined by incorporating a multiplicative pre-factor.
  • Sub-logarithmic Regime: For the slowly growing availability $t=\sqrt{\log_{2}n}$, the empirical decoding failure probability does not appear to vanish over the simulated range. Instead, it saturates at a relatively high error floor (approximately $0.2$ to $0.4$), suggesting that sub-logarithmic availability may be insufficient for reliable error correction via majority voting at practical blocklengths. Meanwhile, the theoretical upper bound in this regime remains close to $1$ for the considered blocklengths, but it does decrease with $n$; its decay is much slower (and therefore less apparent in the plot) than in the linear and polylogarithmic regimes.
@@ IV-B Block Decoding Failure Probability
Figure 2 focuses on the upper bound of the BLER, derived via the union bound on the Bit Error Rate (BER) (i.e., $P_{\mathrm{fail,block}}\leq nP_{\mathrm{fail,bit}}$) for a BSC with $p_{f}=0.13$. Since this BLER is derived from a union bound, the resulting values may exceed both the true BLER and even $1$, which is reflected in the vertical axis of the plot. We do not report empirical BLER values, as their computation becomes infeasible for large block lengths: verifying correct block decoding requires checking the correctness of all symbols, each of which involves an increasing number of recovery sets as $n$ grows. This figure provides a fine-grained analysis of the polylogarithmic regime by varying the exponent of the logarithmic growth, specifically $t=(\log_{2}n)^{\alpha}$ for $\alpha\in\{1.8,1.9,2.05\}$, and highlights the sensitivity of block-level performance to the exponent $\alpha$. For smaller values of $\alpha$ (e.g., $\alpha=1.8$), the BLER bound remains relatively large over the range of blocklengths considered and decreases only at very large $n$, indicating that such growth of $t(n)$ is insufficient for practical reliability. In contrast, slightly larger exponents (e.g., $\alpha\geq 1.9$) lead to a rapid decay of the failure probability, illustrating that the availability must scale with a sufficiently large power of $\log n$ to offset the accumulation of error probability as the block length increases.
@@ Acknowledgment
This work was supported in part by the NSF-BSF grant FET212026. The authors thank A. Barg for providing them with the Russian original and its English translation of [17].